Ticket Number: HEA-NOC/20090331-13 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20090331 18:27 UTC+1 Problem Start: 20090401 08:15 UTC+1
Ticket Closed: 20090401 10:50 UTC+1 Problem End: 20090401 09:15 UTC+1
Site/line: IPv6, IPv6 tunnels
Due to previous problems with IPv6 traffic travelling via LINX, London, we
have been asked to assist in a debugging effort to determine what the root
cause of the problem was.
In order to do this, one of our transit suppliers will restore a route for
IPv6 traffic through LINX, and we will test for IPv6 packet loss for flows
routed through that PoP.
There may be an impact to IPv6 traffic routed to the general internet via
LINX. Up to 10% of packets may be lost during the maintenance window,
assuming we replicate the problem that existed before.
No impact to IPv4 services is expected.
20090331 18:18 GAR Created ticket to track work.
20090401 10:48 GAR Debugging information was captured, and customers may
have noticed approximately 10% packet loss of IPv6 traffic only during the
15 minutes while a known defective configuration was applied for
investigative purposes. IPv4 traffic was unaffected.
Traffic is now routed via Tiscali to DCIX, rather than through LINX, and
we are not seeing any IPv6 packet loss on test routes. Closing ticket.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20090331-13
HEAnet Limited. Registered in Ireland, No. 275301.